The javascript on this page is used to preload the images for links and to swap the images used for links
go to content
layout image go to introduction
go to science datago to mission status reportgo to location / positiongo to 3d data mapsgo to moon VRMLgo to further reading
layout image
Layout imagego to home
go to projectgo to resultsGo to datavisgo to historyGo to scienceGo to educationGo to archivesgo to resources


Lunar Prospector Status Report #18

February 9, 1998 - 12:00 noon EST (3:00 p.m. PST)

The Lunar Prospector Mission Control Center at NASA's Ames Research Center reported that, as of 8:21 p.m. (PST) on Sat., Jan. 31, 1998, the spacecraft continues to operate well. The past weekend, a minor glitch with the spectrometer instrument electronics occurred, in which the incoming data stream for the vehicle's three spectrometers was temporarily halted. Mission controllers decided to power the spectrometer electronics subsystem on and off, which corrected the problem. All instruments were reconfigured and reported to be operating properly. A total of 38 commands were communicated to the spacecraft to resolve this situation.

Current statistics regarding the spacecraft and its instruments, as reported by Mission Operations Manager Marcie Smith (at 10:00 a.m. (PST) on Feb. 2, 1998) are as follows:

Orbit Number: 256
Data Downlink Rate: 3600 bps
Spacecraft Spin Rate: 11.94 rpm
Spin Axis Attitude: Longitude: 251 deg
Latitude: 89.2 deg
Trajectory: Periselene: 96 km
Aposelene:
104 km
Period: 118 minutes duration
Occultations: 46 minutes duration
Eclipses: 32 minutes duration